The pregnancy length in days for a population of new mothers can be approximated by a normal distribution with a mean of 264 days and a standard deviation of 8 days. ​(a) What is the minimum pregnancy length that can be in the top 9​% of pregnancy​ lengths? ​(b) What is the maximum pregnancy length that can be in the bottom 5​% of pregnancy​ lengths? ​(a) The minimum pregnancy length is days? ​(Round to one decimal place as​ needed.)
Answer #1

solution:

Given that,

mean = \mu = 264

standard deviation = \sigma =8

Using standard normal table,

P(Z > z) = 9%

= 1 - P(Z < z) = 0.09

= P(Z < z ) = 1 - 0.09

= P(Z < z ) = 0.91  

z =1.34 (using standard normal (Z) table )

Using z-score formula  

x = z * \sigma + \mu

x= 1.34 *8+264

x= 274.72

b.

Given that,

mean = \mu = 264

standard deviation = \sigma = 8

Using standard normal table,

P(Z < z) = 5%

= P(Z < z) = 0.05

= P(Z <-1.65 ) = 0.05

z =-1.65 Using standard normal table,

Using z-score formula  

x= z * \sigma + \mu

x= -1.65 *8+264

x= 250.8